OpenStudy (anonymous):

Find the exact area of a circle having the given circumference. 8 pi

OpenStudy (amistre64):

in order to find the area we need to integrate with respect to the radius :)

OpenStudy (amistre64):

circumference = 2pi r 8pi = 2pi r r = 8pi/2pi = 4 \[\int_{0}^{4}(2\pi\ r)dr\]

OpenStudy (bahrom7893):

dang lol amistre she doesn't know integrals yet

OpenStudy (bahrom7893):

Solve for the radius and then use A=pi r^2

OpenStudy (amistre64):

\[2pi\int_{0}^{4}r\ dr\] \[2pi(\frac{r^2}{2})|^{4}\] \[pi (4^2)=16pi\]

OpenStudy (bahrom7893):

oh sorry nevermind u already solved for r, didn't see that

OpenStudy (bahrom7893):

No wait she doesn't know integrals yet

OpenStudy (amistre64):

lol it aint forbidden knowledge ;)